how would I factor the equation (c + 3)^2 - 49
Found 2 solutions by jainenderkapoor, rapaljer:
Answer by jainenderkapoor(61)

The question can be factorised in the following manner :-
(c + 3)^2 - 49
(c + 3)^2 - 7^2
Now using the formula that a^2 - b^2 = (a + b) (a - b)
(c + 3 + 7) (c + 3 - 7)
So we get (c + 3 + 7) = 0
c + 10 = 0
c = -10
(c + 3 - 7) = 0
c - 4 = 0
c = 4
Hope this is clear to you.
In case you want to learn more about the topic, please do write to me at kapoorjai1@rediffmail.com
I am avaliable for online tutoring.


Answer by rapaljer(4551)

This is NOT an equation, so you can't solve for c, as was done in the previous solution that was posted by a tutor. The problem is to FACTOR the expression, NOT to SOLVE the equation!!
It should be
( (c+3) - 7 )( (c+3) + 7)

This simplifies to
(c-4)(c+10)

R^2 at SCC